OpenStudy (anonymous):

rationalize the denominator assume that all expressions under radicals represent positive numbers

OpenStudy (anonymous):

\[\sqrt{u}-\sqrt{v}/\sqrt{u}+\sqrt{v } \]

OpenStudy (anonymous):

Multiply top and bottom by sqrt u - sqrt v

OpenStudy (anonymous):

im not sure what you mean

OpenStudy (anonymous):

please help im not getting this one

OpenStudy (anonymous):

If u multiply the bottom by sqrt u - sqrt v it becomes (u-v) so now multiply the top by the same sqrt u - sqrt v